### v1

# Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs

Abstract
Solving Partially Observable Markov Decision Processes (POMDPs) in continuous state, action and observation spaces is key for autonomous planning in many real-world mobility and robotics applications. Current approaches are mostly sample based, and cannot hope to reach near-optimal solutions in reasonable time. We propose two complementary theoretical contributions. First, we formulate a novel Multiple Importance Sampling (MIS) tree for value estimation, that allows to share value information between sibling action branches. The novel MIS tree supports action updates during search time, such as gradient-based updates. Second, we propose a novel methodology to compute value gradients with online sampling based on transition likelihoods. It is applicable to MDPs, and we extend it to POMDPs via particle beliefs with the application of the propagated belief trick. The gradient estimator is computed in practice using the MIS tree with efficient Monte Carlo sampling. These two parts are combined into a new planning algorithm Action Gradient Monte Carlo Tree Search (AGMCTS). We demonstrate in a simulated environment its applicability, advantages over continuous online POMDP solvers that rely solely on sampling, and we discuss further implications.
1 Introduction
Figure 1: Action adaptive MIS tree with two levels of actions and children state nodes. Usual states for MDPs are replaced by belief states for POMDPs. The actions in the action nodes are updated as the search progresses, indicated with a darker shade, and new states are sampled from the newest actions, denoted in the same color. The MCTS selection path, highlighted in blue, chooses at action nodes the latest action, and then proceeds at random to a child state.
The Markov Decision Process (MDP) and the Partially Observable Markov Decision Process (POMDP) are flexible mathematical frameworks for modeling decision-making under uncertainty. Computing policies for (PO)MDPs that are in continuous in states, actions and observations, has mostly focused on global search algorithms based on POMCP [?], like POMCPOW and PFT-DPW [?], that explore the action space only by sampling. Several algorithms have been suggested for mixed continuous settings, in which not all the state, action and observation spaces are continuous. For continuous state and action spaces, GPS-ABT [?] and LCEOPT [?] use generalized pattern search and cross-entropy optimization respectively, both of which are gradient-free optimization methods. For continuous state and observation spaces, AdaOPS [?] suggested belief packing to combine values for similar beliefs, and LABECOP [?] introduced lazy belief extraction to increase belief diversity.
The first example of an algorithm with performance guarantees for continuous POMDPs was VOWSS [?], that combined Voronoi Optimistic Optimization [?] for action sampling with Weighted Sparse Sampling [?]. The authors also suggested a practical version of the algorithm, VOMCPOW, that combines MCTS with Double Progressive Widening (DPW). [?] presented hierarchical Voronoi-based action sampling, showing improved performance at action dimensionality of 6 to 12.
Our research is motivated by the idea that in many continuous settings, gradient-based optimization is often much faster to converge than gradient-free optimization. It might be prohibitively expensive to sample enough actions to statistically find a near-optimal one. In our approach we consider non-parametric beliefs, making it applicable to a wide range of problems, specifically not restricted by Gaussian or other parametric assumptions. Our methodology consists of two parts. First, we consider how to update actions during search time, without losing information of previous action-value estimates. Second, we derive value gradients w.r.t. actions, for the purpose of computing action gradient updates. We hypothesize that incorporating gradient optimization into continuous (PO)MDP planning will explore the action space more efficiently, especially at high dimensional settings.
A similar idea has been recently explored in the algorithm VG-UCT by [?], where the local gradient-based optimization is combined with the global-search of sampled-based MCTS, but with several drawbacks compared to our approach. Their gradient computation approach is computationally expensive due to being based on finite differences and being recursive. They assume a deterministic transition model given a noise parameter, which is problematic for deriving gradients in POMDPs. Lastly, they do not address the distribution shift that occurs from action updates, which affect the action-value estimates.
Our approach requires access to the state transition model’s likelihood. This is in contrast to previous online (PO)MDP solvers and model-free approaches that consider only use of a generative model. Yet, recent advances in physical simulators have included scalable differentiable physics simulators [??], and GPU-parallelization that have allowed generating accurate simulations with speedups in the scale of orders of magnitudes [??]. We foresee that this will allow to compute the differentiable physical transition models more easily in the near future.
1.1 Additional Related Work
Information sharing in discrete MDPs with repeating states was considered in [?]. In POMDPs, previous knowledge utilization via Multiple Importance Sampling (MIS) [?] was introduced in [?]. We share information only between sibling action branches, which eliminates a costly search through a belief database. MIS in POMDP planning for sample reuse was also considered in [?], however in the context of calculating observation likelihoods.
KR-UCT [?] use kernel regression to estimate the value of actions by sharing information between sibling action branches, and for the proposal of new actions. They consider MDPs only, and do not consider updating existing actions. In similar vein, Bayesian optimization for information sharing between different actions was explored in kCBTS [?] and BOMCP [?]. These methods are purely heuristic, and the information sharing is used only for the sampling of new actions, whereas we consider updating existing actions.
In [?], BetaZero was suggested to share information by learning offline policy and value networks over parametric beliefs. However, they consider discrete action spaces, they do not share information between action branches online, nor do they handle modeling inaccuracy.
Our gradient results can be seen as similar to the policy gradient theorem [?] or to the deterministic policy gradient theorem [?], where we calculate gradients w.r.t. actions, rather than policy parameters.
1.2 Contributions
Our work is the first to consider an action-gradient of the value function based on transition likelihoods in continuous (PO)MDPs. Specifically, our main contributions are:
  1. 1.
    We formulate a novel search tree with recursive MIS estimators, for updating actions with consistent value estimates in continuous (PO)MDPs. We show efficient update equations for the self-normalized MIS heuristic.
  2. 2.
    We provide a novel derivation for action gradients of the value function in (PO)MDPs based on transition likelihoods. This gradient computation can be computed at a single action branch at a time, unlike previous MDP value gradients. The extension to POMDPs makes this the first derivation of POMDP value gradients with belief-dependent policies to the best of our knowledge.
  3. 3.
    We derive a novel algorithm Action Gradient MCTS (AGMCTS) that combines the MIS tree with action gradients. This algorithm is applicable to both MDPs and POMDPs. We show several algorithmic improvements and heuristics to make it more efficient.
  4. 4.
    We compare AGMCTS to PFT-DPW and POMCPOW in a simulated POMDP scenario at several dimensions and different noise levels in the domain knowledge. AGMCTS had the best overall performance on a per-simulation budget.
2 Background
MDPs.
An MDP is the tuple $\langle\mathcal{S},\mathcal{A},p_{T},r,\gamma,L,b_{0}\rangle$. $\mathcal{S},\mathcal{A}$ are the state and action spaces. The transition model $p_{T}(s^{\prime}\mid s,a)$ models the probability of arriving at state $s^{\prime}$ when taking action $a\in\mathcal{A}$ at state $s\in\mathcal{S}$. The reward function $r(s,a,s^{\prime})\in\mathbb{R}$ gives the immediate reward of transitioning from state $s$ to $s^{\prime}$ by taking action $a$. We define the reward of $(s,a)$ as the expected reward over outcome states $r(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$. $\gamma\in(0,1]$ is the discount factor. The MDP starts at time 00 and terminates after $L\in\mathbb{N}\cup\{\infty\}$ steps. The initial state is drawn from the distribution $s_{0}\sim b_{0}$.
A timed policy at time $i$, denoted as $\pi_{i}$, is a mapping $\pi_{i}\colon\mathcal{S}\to\mathcal{A}$. A policy $\pi$ is a collection of timed policies from the start time until the horizon: $\pi\triangleq(\pi_{t})_{t=1}^{t=L}$. The value function of a policy $\pi$ at time $t$ is the expected sum of discounted rewards until the horizon: $V_{t}^{\pi}(s_{t})\triangleq\mathbb{E}_{s_{t+1:L}\mid s_{t},\pi}[\sum_{i=t}^{L}\gamma^{i-t}r(s_{i},\pi_{i}(s_{i}),s_{i+1})]$. The action-value function is often used as an intermediate step in the estimation of the value function, and is defined as $Q_{t}^{\pi}(s,a)\triangleq\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime})]$, where state $s^{\prime}$ is at the proceeding time of $s$. The goal of planning is to compute a policy that maximizes the value function.
POMDPs.
A POMDP adds $\langle\mathcal{O},p_{O}\rangle$ to the MDP tuple. The observation space is $\mathcal{O}$, and the observation model $p_{O}(o\mid s)$ is the conditional probability of receiving an observation $o\in\mathcal{O}$ at $s\in\mathcal{S}$. A history at time $t$, $H_{t}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1},o_{t})$, is defined as a sequence of the starting belief, followed by actions taken and observations received until $t$. Because of the partial observability, the agent has to maintain a probability distribution of the current state given past actions and observations, known as the belief. The belief space, denoted in $\mathcal{B}$, is the space of all probability distributions over $\mathcal{S}$. The belief at time $t$ is defined as the conditional distribution of the state $b_{t}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t})$, given the history. We denote $H_{t}^{-}\triangleq(b_{0},a_{0},o_{1},\dots,a_{t-1})$ for the same history without the last measurement, and the propagated belief $b_{t}^{-}(s_{t})\triangleq\mathbb{P}(s_{t}\mid H_{t}^{-})$. It can be shown that optimal decision-making can be made given the belief, instead of considering the entire history [?]. Since the belief fulfills the Markov property, a POMDP is an MDP on the belief space, commonly referred to as belief MDP. Policies, the value function and the action-value function, extend from MDPs to POMDPs by equivalent definitions on $b\in\mathcal{B}$ as states. As an example, $r(b,a)\triangleq\mathbb{E}_{b^{\prime}\mid b,a}[r(b,a,b^{\prime})]$. In the general case, the belief-reward function may be any function of beliefs, f.e. entropy, and this is called $\rho$-POMDP [?]. Often, the belief reward has an underlying state-reward, in the form $r(b,a,b^{\prime})=\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]$. We refer to this case as a state-reward POMDP.
Importance Sampling (IS).
The Monte-Carlo (MC) estimator of an expectation, denoted with $\hat{\mathbb{E}}$ or f.e. $\hat{V}_{t}^{\pi}$, is its approximation with a finite number of samples. Importance Sampling (IS) is a technique in MC estimation where a proposal distribution $q$ is used to sample from instead of $p$. The resulting estimator for $f(x)$ is $\hat{f}=\frac{1}{N}\sum_{i=1}^{N}\frac{p(x^{i})}{q(x^{i})}f(x^{i})$. This estimator is consistent if $q(x)=0\Rightarrow p(x)=0$. With a proper choice of $q$ the variance of the estimator can be reduced, or we can calculate estimates for a target $p$ which is difficult to sample directly. Commonly the sampled weights $w_{i}=p(x^{i})/q(x^{i})$ are normalized to sum to $1$, giving the self-normalized IS estimator [?, 1.3.2].
Particle Beliefs.
Often in continuous settings we cannot maintain an exact belief. Hence, particle filters are used instead to represent a belief non parametrically [?]. We denote particle beliefs with $\bar{b}$. The particle belief of $J$ particles is represented by a set of state-weight pairs $\bar{b}=\{(s^{j},w^{j})\}_{j=1}^{J}$, and is defined as the discrete distribution $\mathbb{P}(s\mid\bar{b})={\sum_{j=1}^{J}{w^{j}\cdotp\delta(s-s^{j})}}/{\sum_{j=1}^{J}{w^{j}}}$. The basic particle filtering algorithm, Sequential Importance Sampling (SIS) [?, 1.3.2], recursively updates and normalizes the importance weights of the particles based on the observation likelihood. Sequential Importance Resampling (SIR) adds resampling of state particles after the weight updates, to avoid weight degeneracy [?]. The particle belief MDP (PB-MDP) was considered in [?], where the belief space is explicitly the space of ordered particle beliefs $\bar{b}$ of $J$ particles. An ordered particle belief is the ordered set $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$.
Multiple Importance Sampling (MIS).
MIS [?] is a method to combine the advantages of several sampling methods. It is defined as $F=\sum_{i=1}^{n}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}w_{i}(x^{i,j})\frac{f(x^{i,j})}{q_{i}(x^{i,j})}$, which is a weighted average of $n$ IS estimators. If the weights $w_{i}$ satisfy $\sum_{i=1}^{n}w_{i}(x)=1$ whenever $f(x)\neq 0$, and $w_{i}(x)=0$ whenever $q_{i}(x)=0$, then $F$ is unbiased. The choice of $w_{i}$ may depend on the samples $x^{i,j}$. The weighting strategy $w_{i}(x)=\frac{n_{i}q_{i}(x)}{\sum_{k}n_{k}q_{k}(x)}$, named the balance heuristic, has its variance provably bounded from the optimal weighting strategy.
Monte Carlo Tree Search (MCTS).
MCTS is an algorithm used to quickly explore large state spaces [?]. It iteratively repeats four steps to build a search tree that approximates the action-values, using a best-first strategy: (1) Selection Starting from the root node, descend recursively until a node to which children can be added is found; (2) Expansion A new node is added as a child, according to action expansion strategy; (3) Simulation A simulation is run from the new node according to the rollout (default) policy; (4) Backpropagation The simulation result is backpropagated through the selected nodes to update the action-values. Often UCT is used at selection to balance between exploration of new actions and exploitation of promising ones [?]. Double Progressive Widening (DPW) [?] is a technique to limit the branching factor from being infinite in continuous settings. The number of children of a node is artificially limited to $kN^{\alpha}$ for $N$ visitations, for hyperparameters $k$ and $\alpha$. The progressive widening is double due to being applied both to the branching of actions, and branching of states.
Stochastic Gradient Optimization (SGO).
SGO is a widely used technique for efficient optimization of differentiable functions [?]. Gradient optimizers tremendously rose in popularity to their applications in deep neural networks. Adam [?] is a widely used algorithm, with heuristically adaptive step size based on first and second moment estimates of the gradient. Adam has been shown to perform well with noisy gradient estimates, and requires minimal tuning to different problems.
3 Methodology
3.1 Action-Values with Importance Sampling
Figure 2: Recursive definitions of the MIS tree, for the particle belief MDP (PB-MDP) case, in which states are particle beliefs. State nodes are circles and action nodes are squares. The state $\bar{b}$ has a value estimate $\hat{V}(\bar{b})$ and visitation count $n(\bar{b})$. The action node $(\bar{b},a)$ has an immediate reward estimate $\hat{r}(\bar{b},a)$, a future value estimate $\hat{V}_{f}(\bar{b},a)$, and visitation count $n(\bar{b},a)$. The action-value estimate is defined via $\hat{Q}(\bar{b},a)\triangleq\hat{r}(\bar{b},a)+\gamma\hat{V}_{f}(\bar{b},a)$. During action update we update $(\bar{b},a)$ to $(\bar{b},a^{\prime})$. We highlight that in the PB-MDP, we also store the propagated belief $\bar{b}^{-}$. The child state node has all properties as $\bar{b}$, and stores the proposal action $a_{prop}(\bar{b}^{-})$. During action update (11) we do $\omega^{\prime}={p_{T}(\bar{b}^{-}\mid\bar{b},a^{\prime})}({p_{T}(\bar{b}^{-}\mid\bar{b},a_{prop}(\bar{b}^{-}))})^{-1}$. In this example, $a_{prop}(\bar{b}^{-})=a$.
In an MDP setting, when updating action $a$ to $a^{\prime}$, we would like to estimate the new action-value $Q_{t}^{\pi}(s,a^{\prime})$ by reusing the information that was used for estimating $Q_{t}^{\pi}(s,a)$. Otherwise, updating actions would require us to theoretically recompute the entire subtree, which would render the algorithm very inefficient. Since this process might occur several times for each action branch, in the general case each child state may have been generated from a different proposal action. Hence, we have to account for an MIS scenario in order to estimate $Q_{t}^{\pi}(s,a^{\prime})$.
Assuming that $p_{T}(s^{\prime}\mid s,a^{\prime})$ is a non-strict subset of the support of $p_{T}(s^{\prime}\mid s,a)$, we can formally write
$\textstyle Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))].$ (1)
For POMDPs, the posterior belief transition likelihood is generally intractable. However, by using the law of total expectation, we introduce a dependence of the posterior belief $b^{\prime}$ on the propagated $b^{-}$. This makes only $b^{-}$ directly dependent on $a$, which allows us to introduce an importance ratio over the propagated belief rather than posterior:
$\textstyle Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\textstyle[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))].$ (2)
This is known as the propagated belief trick [?]. We can see that (1) and (2) are analogous, except for the importance ratio being between propagated beliefs in the POMDP case. The propagated belief likelihood is rather simple to calculate for appropriate choice of the belief space.
Lemma 1.
Let $\bar{b}^{-}=((s^{-,j},w^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$ by SIS or SIR. Hence, the $j$th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and:
$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a).$ (3)
We refer to the appendix for all proofs
3.2 Action Adaptive MIS Trees
We now present our definitions of the novel MIS tree, a recursive tree structure with MIS value estimates, that allows sharing value information between sibling actions. Similarly to belief trees, the tree has levels alternating with states and action nodes, starting from the root state node. Yet in contrast, the MIS tree supports replacing an action node with a new action, and updating the value estimates of the new action and its parent state node based on previous samples. We later introduce action updates via gradient-based optimization.
For the following discussion, we present notations for recursively defining estimators at a state node $s$ in the MIS tree, or some action node $(s,a)$. The generalization to POMDPs is done by taking ordered particle belief states $\bar{b}$, and the following definitions are analogous, illustrated in Figure 2.
Denote the children actions of $s$ as $C(s)=\{a^{i}\}_{i=1}^{\lvert{C(s)}\rvert}$, and the children states of $(s,a)$ as $C(s,a)=\{s^{\prime,i}\}_{i=1}^{M}$, where $M\triangleq\lvert{C(s,a)}\rvert$. The visitation counts of state and action nodes are denoted $n(s)$ and $n(s,a)$ respectively. We denote the proposal action $a_{prop}(s)$, the action that was used to sample $s$, as possibly $a_{prop}(s^{\prime})\neq a$ for $s^{\prime}\in C(s,a)$. We assign each action node $(s,a)$ an action-value estimate $\hat{Q}(s,a)$ that estimates $Q_{t}^{\pi}(s,a)$, assuming $s$ is a state at time $t$. Compared to traditional MCTS algorithms, we also estimate the value function at each state node. We define it as a weighted mean:
$\textstyle\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}).$ (4)
The following two relationships between visitation counts are maintained in MCTS:
$\textstyle n(s)=\sum_{C(s)}n(s,a^{i}),\;\textstyle n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (5)
where $n(s^{\prime,i})_{+1}\triangleq n(s^{\prime,i})+1$. The asymmetry is the result of initializing state nodes to $n(s)=0$ when assigned a rollout value, whereas action nodes are initialized to $n(s,a)=1$. This allows value estimates to be updated completely when they are visited again, and for their parent’s action-value estimate to remember the initial rollout estimate.
We split $\hat{Q}(s,a)$ into two estimators: (1) Immediate reward estimate $\hat{r}(s,a)$, for $\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})]$; (2) Future value estimate $\hat{V}_{f}(s,a)$ for $\mathbb{E}_{s^{\prime}\mid s,a}[V^{\pi}(s^{\prime})]$ of the proceeding time, such that if $s$ is a state at time $t$, then $\hat{V}_{f}(s,a)$ estimates the expected $V_{t+1}^{\pi}(s^{\prime})$. Together they define $\hat{Q}(s,a)\triangleq\hat{r}(s,a)+\gamma\hat{V}_{f}(s,a)$.
Generally in MIS, each proposal distribution $q_{i}$ has $n_{i}$ samples. In our situation, we assume that each of $C(s,a)$ may have been sampled from a different action, has its own visitation counter and an associated single value estimate $V(s^{\prime,i})$. Therefore, we adapt the most naive unbiased weighting, that is to give equal weight to all samples, and weight the posterior states by the visitation counts. Hence, the naive MIS estimator for the future value is
$\textstyle\hat{V}_{f}(s,a)=(n(s,a))^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (6)
where $\omega^{i}={p_{T}(s^{\prime,i}\mid s,a)}({p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i}))})^{-1}$. The naive MIS estimator for $\hat{r}(s,a)$ is identical, with $\hat{V}(s^{\prime,i})$ replaced by $r(s,a,s^{\prime,i})$.
In practice, this estimator has very bad performance. On the other hand, the balance heuristic requires $O(M)$ time to update [?], which can be quite expensive. We found a simple solution which is to self-normalize the weights. In regular IS, it is known to reduce the variance of the estimator at the expense of introducing bias, and asymptotically converges to the true estimate under weak assumptions [?, 1.3.2]. We define self-normalized MIS (SN-MIS) by
$\textstyle\eta\triangleq\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1},$ (7)
$\textstyle\hat{V}_{f}(s,a)=\eta^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (8)
and again the SN-MIS estimator for $\hat{r}(s,a)$ is defined by replacing $\hat{V}(s^{\prime,i})$ with $r(s,a,s^{\prime,i})$.
3.3 Updating MIS Trees in MCTS
We now discuss how to update the SN-MIS estimators in the steps of the MCTS routine. We highlight that for any other MIS weighting heuristic, a similar analysis would follow.
Currently, we do not address selection, and we assume the UCT criterion is not altered. Simulation is unaffected, as it is determined only by the rollout policy. Hence, we analyze the different subcases of backpropagation and expansion steps. We add a new step to MCTS named action update, where we update action node $(s,a)$ to $(s,a^{\prime})$ with the new action $a^{\prime}$, and we show its update equations. In general, action update can be applied at every point of MCTS.
For the recursive analysis, we assume backpropagated values for the children nodes are already updated, and at each step we only update an action node $(s,a)$ and its parent node $s$. The updated values are denoted with a prime, and we assume that we have access to the values before the update. The transition likelihood terms $w_{i}$ are usually the most expensive to calculate, hence it is important to cache them, and recalculate the numerator of $w_{i}$ only when necessary. The derivations of the update equations, alongside numerically stable forms with log-likelihoods are given in the appendix.
Action Backpropagation.
Let $s^{\prime,i}$ be an updated node. Hence, we updated $n(s^{\prime,i})$ to $n^{\prime}(s^{\prime,i})$ and $\hat{V}(s^{\prime,i})$ to $\hat{V}^{\prime}(s^{\prime,i})$. We update $n(s,a)$ by (5) and perform
$\textstyle\eta^{\prime}=\eta+\omega^{i}(n{{}^{\prime}}(s^{\prime,i})-n(s^{\prime,i})),$ (9)
$\textstyle\hat{V}_{f}^{\prime}(s,a)=(\eta^{\prime})^{-1}(\eta\hat{V}_{f}(s,a)+$
$\textstyle\omega^{i}(n{{}^{\prime}}(s^{\prime,i})\hat{V}^{\prime}(s^{\prime,i})-n(s^{\prime,i})\hat{V}(s^{\prime,i})).$ (10)
State Expansion.
This case is the same as action backpropagation, where the updated index is a new summand at index $i=M+1$. The same equations apply with $n(s^{\prime,i})=0$.
Action Update.
Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega^{i}$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
$\textstyle(\omega^{i})^{\prime}=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))^{-1},$ (11)
and recalculate (7) and (8) with the new ($\omega^{i})^{\prime}$, which are also $O(M)$ time operations. For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
Terminal State Backpropagation.
Let $s$ be a terminal node. Its value estimate is based only on rollouts. Hence, for a new rollout value $v^{\prime}$, we perform a running average update
$\textstyle\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (12)
where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1).
Non-Terminal State Backpropagation.
Let $(s,a^{i})$ be an updated node. Hence, we updated $n(s,a^{i})$ to $n^{\prime}(s,a^{i})$ and $\hat{Q}(s,a^{i})$ to $\hat{Q}^{\prime}(s,a^{i})$. We update $n^{\prime}(s)$ by (5) and perform
$\textstyle\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}(n(s)\hat{V}(s)+n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})$
$\textstyle{}-n(s,a^{i})\hat{Q}(s,a^{i})).$ (13)
3.4 MDP and POMDP Value Gradient
In the previous subsections we’ve assumed that the updated action $a^{\prime}$ is arbitrary. We now present a method to compute an estimate of $\nabla_{a}Q(s,a)$, in order to compute a gradient-based update. Our derivation of $\nabla_{a}Q(s,a)$ is based on the gradient of the transition likelihood, and there are two main advantages to this approach. First, the non-recursive computation can be computed separately for each action branch in the MIS tree, which greatly reduces computational complexity. Second, which is our key observation, is that this derivation extends to POMDPs as well via the propagated belief trick, and has a computable form for the PB-MDP.
Theorem 1.
For well-behaved $p_{T}$, $r$, the MDP action gradient of the action-value function is given by the expectation
$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})\textstyle=\mathbb{E}_{s^{\prime}\mid s,a}[\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
$\textstyle\cdot(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime}))].$ (14)
Additionally, for well-behaved $p_{T}$, $p_{O}$, $r$, the POMDP action gradient of the action-value function can be computed by replacing posterior belief likelihood terms with propagated belief likelihood terms, i.e. $b_{t+1}^{-}$ instead of $b_{t+1}$ in all $p_{T}$ terms.
The gradient is written such that states $s^{\prime}$ could have been sampled from an action $a$ different from $a^{\prime}$. Intuitively, it tends toward actions that both maximize immediate rewards, and maximize the likelihood of future states with high values. Importantly, because $s^{\prime}$ is an integration variable, the future value is not a function of $a^{\prime}$, and has no gradient w.r.t. it.
Theorem 1 is general, and applies to MDPs, state-reward POMDPs, and $\rho$-POMDPs. The following Theorem 2 modifies the gradient for state-reward POMDPs. The gradient for the immediate reward is based on single state samples rather than sampling full beliefs. It may give more accurate gradient estimates, albeit possibly more expensive to compute.
Theorem 2.
For a state-reward POMDP, the action gradient of the action-value function is given by:
$\textstyle\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a)=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})$
$\textstyle+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]\textstyle+\gamma\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}[\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$\textstyle\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})].$ (15)
For ordered particle beliefs, the gradient log of the propagated transition likelihood from Lemma 1 is given by:
$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)=\sum_{i=1}^{J}\nabla_{a}\log p_{T}(s^{-,j}\mid s^{j},a).$ (16)
3.5 Action Gradient Monte Carlo Tree Search
Algorithm 1 AGMCTS
procedure S IMULATE$(s,d)$
1:  if
I
S
T
ERMINAL$(s)$ OR $d=0$ then
2:     $v\leftarrow$
R
OLLOUT$(s)$Rollout until terminal state}
3:      U PDATE T ERMINAL($s,v$) {Eq. (12)}
4:     return $v$
5:  $a\leftarrow$
A
CTION
P
ROG
W
IDEN$(s)$
6:   $addSample\leftarrow$ A CTION O PT $(s,a,d)$
7:  if  $\lvert{C(sa)}\rvert\leq k_{o}N(sa)^{\alpha_{o}}$ OR $addSample$  then
8:     $s^{\prime},r\leftarrow G(s,a)$Replace $s^{\prime}$ by $(b^{\prime},b^{-})$ for POMDP}
9:     $C(sa)\leftarrow C(sa)\cup\left\{(s^{\prime},r)\right\}$
10:     v
R
OLLOUT
(s,d1)
[…]
R
OLLOUT
[…]
11:  else
12:     $s^{\prime},r\leftarrow$ sample uniformly from $C(sa)$
13:     v
S
IMULATE
(s,d1)
[…]
S
IMULATE
[…]
14:   U PDATE MIS ($s,a,s^{\prime},r$) {Eqs. (9), (10), (13)}
procedure
A
CTION
O
PT
(s,a,d)
A
CTION
O
PT
[…]
1:  $addSample\leftarrow$ FALSE
2:  for all $k=1,\dots,K_{opt}$ do
3:     $g_{a}^{q}=\hat{\nabla}_{a}\hat{Q}(s,a)$Eq. (19) or (17)}
4:     Initialize $a^{acc}\leftarrow a$ for $(s,a)$ if required
5:     aacc
O
PT
(s,aacc,gaq)
[…]
O
PT
[…]
{Adam/other optimizer}
6:     if $\lVert{a^{acc}-a}\rVert>T_{d_{a}}$ then
7:        $addSample\;|\mathbin{=}\;$
A
CTION
U
PDATE$(s,a,a^{acc})$Eqs. (11), (13); Return TRUE if all $\omega^{i}\leq T_{\omega}^{add}$.}
8:  return $addSample$
We introduce Action Gradient MCTS (AGMCTS), a novel online planning algorithm based on MCTS with DPW. It is applicable to POMDPs via ordered particle belief states, like PFT-DPW [?]. It is described in high-level in Algorithm 1, with novel components highlighted in blue. It combines the MIS estimators for the action-value function, with gradient-based action optimization, and updates actions accordingly. AGMCTS applies action update before expansion. This allows us to pass the information from to determine whether to expand the current action node.
Monte Carlo Gradient Estimation.
While computing (14), (15) can be done based on all children of the current action node, we employ the idea of stochastic gradient optimization to MC estimate the gradient. The first estimate is to sample $K_{O}^{\nabla}$ observation branches with weights $\omega^{i}$. It is possible to uniformly sample and weight by $\omega^{i}$, but we found it to be less effective. The MC approximation of (14) becomes
^$\textstyle\hat{\nabla}_{a}\hat{Q}(s,a)=\frac{1}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}\nabla_{a}\log p_{T}(s^{\prime,i_{k}}\mid s,a)$
$\textstyle\cdot(r(s,a,s^{\prime,i_{k}})+\gamma\hat{V}(s^{\prime,i_{k}}))+\nabla_{a}r(s,a,s^{\prime,i_{k}}).$ (17)
For POMDPs, we estimate (16) with $K_{b}^{\nabla}$ state particles via
$\textstyle\nabla_{a}\log p_{T}(\bar{b}^{-}\mid\bar{b},a)\approx\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}\nabla_{a}\log p_{T}(s^{-,j_{l}}\mid s^{j_{l}},a),$ (18)
where indices $j_{l}$ are sampled uniformly from $[1,\dots,J]$. In state-reward POMDPs, we estimate the immediate reward separately. We generate $K_{r}^{\nabla}$ posterior states by sampling $s_{new}^{j}\sim p_{T}(\cdot\mid\bar{b},a)$. The MC estimate of (15) is then
^$\textstyle\hat{\nabla}_{a}\hat{Q}(\bar{b},a)=\frac{1}{K_{r}^{\nabla}}\big{(}\sum_{i=j}^{K_{r}^{\nabla}}\nabla_{a}\log p_{T}(s_{new}^{j}\mid s,a)\cdot$
$\textstyle r(s,a,s_{new}^{j})+\nabla_{a}r(s,a,s_{new}^{j})\big{)}+\big{(}\frac{\gamma}{K_{O}^{\nabla}}\sum_{k=1}^{K_{O}^{\nabla}}(\frac{J}{K_{b}^{\nabla}}\sum_{l=1}^{K_{b}^{\nabla}}$
$\textstyle\nabla_{a}\log p_{T}(s^{-,i_{k},j_{l}}\mid s^{i_{k},j_{l}},a))\textstyle\cdot\hat{V}(\bar{b}^{\prime,i_{k}})\big{)}.$ (19)
Gradient Optimization.
We’ve found throughout our experiments that the Adam optimizer, utilizing step size normalization, was crucial for consistent behavior due to the high variance of the gradient estimates caused by MC estimation and extremely high variance of the propagated belief transition probability. A slight step size decay was beneficial as well for convergence, allowing to choose larger Adam step sizes. We’ve fixed as a hyperparameter $K_{opt}$ the number of consecutive gradient optimization iterations before simulation. Together with the step size, this controls a trade-off between accuracy and computational complexity.
Lastly, since action updates are significantly more costly than computing the gradient estimates, we set a threshold action distance for action update. After each gradient iteration, we update an accumulator of the gradient steps $a_{t}^{acc}$. Only after passing a threshold distance $\lVert{a_{t}^{acc}-a_{t}}\rVert>T_{d_{a}}$, we update $a_{t}\leftarrow a_{t}^{acc}$. This limits the number of action updates, and acts like a convergence criterion.
Thresholds For Adding/Deleting State Nodes.
Although the SN-MIS estimator normalizes the contribution of each posterior node, we would like to ensure that we have relevant samples after several action updates. As a relevance heuristic, we delete child state $s^{\prime,i}$ from $C(s,a)$ if $\omega^{i}<T_{\omega}^{del}$ after action update. Additionally, if $\omega^{i}<T_{\omega}^{add}$ for all $s^{\prime,i}\in C(s,a)$, we force to sample a new posterior node, overriding the progressive widening limitation. This ensures that we do not waste computational resources on states with low contribution to the estimator $\hat{Q}(s,a)$. Due to the possibility of deleting posterior children, care must be taken when updating visitation counters since they do not monotonically increment anymore. For this reason equations (9), (10), (11), (12), (13) are written using previous and updated visitation counts.
4 Experiments
Numerical simulations were carried out to evaluate the performance of AGMCTS in a POMDP setting, compared to state-of-the-art algorithms POMCPOW and PFT-DPW, whose implementations were taken as the standard in the POMDPs.jl framework [?]. AGMCTS was implemented using POMDPs.jl as well. For gradient computations we used automatic differentiation with Enzyme.jl [?]. The analysis and comparison was based on the code of [?], and similarly we used rollout-first-then-random action sampling.
4.1 D-Continuous Light Dark
(a) Middle of scenario, $t=5$.
(b) End of scenario, $t=6$.
Figure 3: AGMCTS at 2-Continuous Light Dark. The agent’s current state is the red dot, the current belief particles in orange, the next state in black, and the next belief particles in gray, the next observation is the green dot. The goal is the blue ring centered at $(0,2.5)$, $b_{0}$ is the red ring centered at $(0,0)$, and the beacon is the green star at $(2.5,0)$. The reward function for the posterior state is the blue-red heatmap drawn in the background. The shaded area is where the observation noise is $\sigma_{O}>5$. The chosen action branch of AGMCTS is drawn as the blue arrows, getting darker in hue with each action update. We can see that the optimized actions point closer to the goal from the current belief’s mean.
This is a continuous extrapolation of the well-known Light-Dark POMDP. In this domain, $\mathcal{S}=\mathcal{O}=\mathbb{R}^{D}$. The agent starts at a random position on the sphere centered at the origin $S^{D-1}_{0.5}$. The goal position is located at $s_{g}=(\boldsymbol{0}_{D-1},2.5)$. A beacon is located to the side at $s_{b}=(2.5,\boldsymbol{0}_{D-1})$. $T=0.2$ is the goal tolerance, and the POMDP terminates if $\lVert{s-s_{g}}\rVert<T$. The horizon is $L=6$ and $\gamma=0.99$. The action space is the ball $B_{1.5}(\boldsymbol{0}_{D})$. The transition model is a simple Gaussian with added noise: $s^{\prime}\sim\mathcal{N}(s+a,(0.025)^{2}\mathbf{I})$. The observations are the agent’s relative position to the beacon, with a Gaussian added noise that increases with distance: $o\sim\mathcal{N}(s-s_{b},(\sigma_{O}(\lVert{s-s_{b}}\rVert))^{2}\mathbf{I})$, with noise function $\sigma_{O}(x)=\min\{15,0.01(x+x^{8})\}$. The reward function is $r(s,a,s^{\prime})=10\cdot g_{1}(d)-2\cdot g_{2}(d)-0.02d^{2}$, where $d\mathbin{=}\lVert{s^{\prime}-s_{g}}\rVert$. $g_{1}$ is a Gaussian centered at 00 with width $\frac{T}{2}$, $g_{2}$ is a Gaussian centered at $5T$ with width $T$. This reward function only depends on the posterior state. The agent needs to carefully target the goal as the reward decreases quickly from the goal, and advancing at random towards it will incur a penalty due to the moat $g_{2}$.
To simulate more difficult domains, we include a noisy rollout policy. The rollout heads directly to $s_{g}$, but the computed action has an added noise sampled from $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$.
4.2 Performance Evaluation
[…]
D $\sigma_{r}$ POMCPOW PFT-DPW AGMCTS
2 0.1 $\boldsymbol{6.62}\pm 0.07$ $5.71\pm 0.08$ $\boldsymbol{6.67}\pm 0.06$
2 0.2 $\boldsymbol{6.23}\pm 0.07$ $4.74,0.08$ $5.96\pm 0.08$
2 0.3 $\boldsymbol{5.61}\pm 0.08$ $4.01\pm 0.09$ $\boldsymbol{5.49}\pm 0.09$
3 0.1 $3.42\pm 0.08$ $3.63\pm 0.08$ $\boldsymbol{5.47}\pm 0.07$
3 0.2 $1.87\pm 0.08$ $1.79\pm 0.08$ $\boldsymbol{3.82}\pm 0.1$
3 0.3 $1.06\pm 0.07$ $0.29\pm 0.06$ $\boldsymbol{1.58}\pm 0.1$
4 0.1 $2.58\pm 0.07$ $2.36\pm 0.07$ $\boldsymbol{4.26},0.08$
4 0.2 $0.53\pm 0.05$ $0.2\pm 0.05$ $\boldsymbol{1.53}\pm 0.09$
4 0.3 $\boldsymbol{-0.36}\pm 0.03$ $-0.86\pm 0.02$ $-1.13\pm 0.04$
Table 1: Performance results for the different algorithms at D-Continuous Light Dark for $D=2,3,4$ dimensions. AGMCTS and PFT-DPW were run for 500 simulations, and POMCPOW for an equivalent time to PFT-DPW. The rollout policy’s standard deviation was $\sigma_{r}=0.1$ during hyperparameter tuning, and then the algorithms were also tested for $\sigma_{r}=0.2,0.3$. The reported results are mean returns over 1000 scenarios $\pm$ standard error. AGMCTS achieved the best performance overall in the different domains.
5 Conclusions
We introduce a novel methodology for considering dynamic action branches in MCTS in continuous (PO)MDP settings. The MIS tree allows action updates while keeping action-value estimates consistent. The new gradient derivation allows computing the action-value gradient in continuous POMDPs, and is supported by the MIS tree for its practical estimation. Our novel planning algorithm AGMCTS incorporates these two novelties together, and with additional algorithmic heuristics, is shown to achieve improved performance on a per-simulation budget. Our experiments suggest that gradient-based action updates may be more effective at higher dimensions, and where domain knowledge is less accurate.
Practical Considerations.
AGMCTS requires a transition likelihood model, both for MIS updates and for the gradient computation. It is application-dependent whether the computational cost of the gradient computation is worth the improvement in performance, compared to using more simulations. In our experiments, AGMCTS had significantly slower planning sessions compared to PFT-DPW. Each algorithm was run in a single thread, yet particle belief algorithms such as AGMCTS and PFT-DPW are highly parallelizable. Furthermore, AGMCTS could be combined with more sophisticated heuristics, like Voronoi-based action sampling and Bayesian optimization methods, which might boost performance. Finally, it is important to note that many algorithmic choices can be tailored depending on the application, like the MIS heuristic and the gradient optimizer, which may significantly affect run times.
Future Work.
The gradient computation and reuse that we’ve shown for POMDPs are based on the propagated belief, which requires full belief trajectories that might be computationally demanding in some domains. We plan to extend our methods efficiently for state trajectories as well. For the action selection we’ve used the standard UCT, however it is not obvious that it remains theoretically correct for the MIS estimator. It is likely that the exploration bonus varies at action updates, and further investigation is required. Finally, AGMCTS requires significant hyperparameter tuning, including DPW parameters and parameters related to action optimization. It is future research to find algorithms that are more general across different domains.
Acknowledgments
This research was partially supported by the Israel Science Foundation (ISF).
Appendix A Proofs and Derivations
A.1 Lemma 1
Lemma 1.
Let $\bar{b}^{-}=((s^{-,j},w^{j}))_{j=1}^{J}$ be generated from $\bar{b}=((s^{j},w^{j}))_{j=1}^{J}$ by SIS or SIR. Hence, the $j$th state particle was generated via $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$, and the weight was copied, i.e. $w^{-,j}=w^{j}$. The ordered particle belief likelihood is then:
$\textstyle p_{T}(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a).$ (20)
Proof.
In the SIS and SIR algorithms, each particle is sampled $s^{-,j}\sim p_{T}(\cdot\mid s^{j},a)$ independently for each $j=1,\ldots,J$. Therefore:
$p(\bar{b}^{-}\mid\bar{b},a)$ (21)
$=p(((s^{-,j},w^{-,j}))_{j=1}^{J}\mid((s^{l},w^{l}))_{l=1}^{J},a),$ (22)
due to the beliefs being ordered, and the independent sampling, we can factorize as:
$=\prod_{i=1}^{J}p(((s^{-,j},w^{-,j}))\mid((s^{l},w^{l}))_{l=1}^{J},a)$ (23)
$=\prod_{i=1}^{J}p(((s^{-,j},w^{-,j}))\mid(s^{j},w^{j}),a).$ (24)
Finally, since $w^{-,j}=w^{j}$, we can eliminate the weights to obtain
$p(\bar{b}^{-}\mid\bar{b},a)=\prod_{i=1}^{J}p_{T}(s^{-,j}\mid s^{j},a),$ (25)
and we denote $p_{T}(\bar{b}^{-}\mid\bar{b},a)\triangleq p(\bar{b}^{-}\mid\bar{b},a)$. ∎
A.2 Derivations of MIS Tree Update Equations
Here we show the derivation of the update equation, and also show their numerically stable forms using log likelihoods for probability values. We denote the regular and the weighted log-sum-exp function as $LSE$, differentiated by having one or two vector inputs, and they are defined as
$LSE(\boldsymbol{x}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\right),$ (26)
$LSE(\boldsymbol{x}_{1,\ldots,n},\boldsymbol{a}_{1,\ldots,n})\triangleq\log\left(\sum_{i=1}^{n}\exp(\boldsymbol{x}_{i})\cdotp\boldsymbol{a}_{i}\right).$ (27)
To handle negative weights, (27) can be defined to compute the log of absolute value and sign separately of the log argument, such that the sign is multiplied back after exponentiation if required. We also denote a vector with brackets []delimited-[][$\cdot$][ ⋅ ].
We remind the visitation counters relationships:
$n(s)=\sum_{C(s)}n(s,a^{i}),\;n(s,a)=\sum_{C(s,a)}n(s^{\prime,i})_{+1},$ (28)
the state value estimator:
$\hat{V}(s)\triangleq\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i}),$ (29)
the self normalized MIS estimator:
$\eta\triangleq\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1},$ (30)
$\hat{V}_{f}(s,a)=\eta^{-1}\sum_{C(s,a)}\omega^{i}n(s^{\prime,i})_{+1}\hat{V}(s^{\prime,i}),$ (31)
and we also remind that we denoted $M\triangleq\lvert{C(s,a)}\rvert$.
We define the following vectors for each a given action node $(s,a)$ or $(\bar{b},a)$:
  1. 1.
    Child state visitation counters.
    MDP:
    $\boldsymbol{n}\triangleq(n(s^{\prime,i})_{+1})_{i=1}^{M},$ (32)
    PB-MDP:
    $\boldsymbol{n}\triangleq(n(\bar{b}^{\prime,i})_{+1})_{i=1}^{M}.$ (33)
  2. 2.
    Target transition likelihoods.
    MDP:
    $\boldsymbol{p}=(p_{T}(s^{\prime,i}\mid s,a))_{i=1}^{M},$ (34)
    PB-MDP:
    $\boldsymbol{p}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a))_{i=1}^{M}.$ (35)
  3. 3.
    Proposal transition likelihoods.
    MDP:
    $\boldsymbol{q}=(p_{T}(s^{\prime,i}\mid s,a_{prop}(s^{\prime,i})))_{i=1}^{M},$ (36)
    PB-MDP:
    $\boldsymbol{q}=(p_{T}(\bar{b}^{-,i}\mid\bar{b},a_{prop}(\bar{b}^{-,i})))_{i=1}^{M}.$ (37)
  4. 4.
    Immediate rewards.
    MDP:
    $\boldsymbol{r}=(r(s,a,s^{\prime,i}))_{i=1}^{M}$ (38)
    PB-MDP:
    $\boldsymbol{r}=(r(\bar{b},a,\bar{b}^{\prime,i}))_{i=1}^{M}.$ (39)
  5. 5.
    Future value estimates.
    MDP:
    $\boldsymbol{v}=(\hat{V}(s^{\prime,i}))_{i=1}^{M}$ (40)
    PB-MDP:
    $\boldsymbol{v}=(\hat{V}(\bar{b}^{\prime,i}))_{i=1}^{M}$ (41)
From here on the PB-MDP is the same up to different notations as the MDP case, and we will show only the latter. Using these notations the immediate reward and future value estimates can be compactly written:
$\eta=\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\oslash\boldsymbol{q}),$ (42)
$\hat{r}(s,a)=\eta^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{r}\oslash\boldsymbol{q}),$ (43)
$\hat{V}_{f}(s,a)=\eta^{-1}\boldsymbol{1}^{T}(\boldsymbol{n}\odot\boldsymbol{p}\odot\boldsymbol{v}\oslash\boldsymbol{q}),$ (44)
[…]
We store the following values:
  1. 1.
    $l(\eta)$.
  2. 2.
    Log likelihoods $l(\boldsymbol{p})$ and $l(\boldsymbol{q})$.
  3. 3.
    $\boldsymbol{n}^{\prime}$ and previous visitation counters $\boldsymbol{n}$.
  4. 4.
    $\boldsymbol{v}^{\prime}$ and previous future value estimates $\boldsymbol{v}$.
The equations for $\hat{r}(s,a)$ are analogous to $\hat{V}_{f}(s,a)$, and we will not show their derivation explicitly.
Action Backpropagation.
Let $s^{\prime,j}$ be an updated node. Hence, we updated $\boldsymbol{n}_{j}$ to $\boldsymbol{n}^{\prime}_{j}$ and $\boldsymbol{v}_{j}$ to $\boldsymbol{v}^{\prime}_{j}$. We update $n(s,a)$ by (28) and perform
$\eta^{\prime}=\eta+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}\left(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}\right),$ (45)
$\hat{V}_{f}^{\prime}(s,a)=\frac{1}{\eta^{\prime}}\left(\eta\hat{V}_{f}(s,a)+\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j})\right).$ (46)
Therefore, the log-form update equations are:
$l(\eta^{\prime})=LSE\left([l(\eta),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})+l(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j})]\right)$ (47)
$\hat{V}_{f}^{\prime}(s,a)=\exp\Big{(}LSE\big{(}[l(\eta),l(\boldsymbol{p}_{j})-l(\boldsymbol{q}_{j})],$
$[\hat{V}_{f}(s,a),\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}]\big{)}-l(\eta^{\prime})\Big{)}.$ (48)
Proof.
We start with the update for $\eta^{\prime}$:
$\eta^{\prime}-\eta$ (49)
$=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (50)
$=\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}-\boldsymbol{n}_{j}),$ (51)
due to all terms being equal except for $\boldsymbol{n}_{j}^{\prime}\neq\boldsymbol{n}_{j}$. Similarly, for $\hat{V}_{f}(s,a)$:
$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta}{\eta^{\prime}}\hat{V}_{f}(s,a)$ (52)
$=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta}{\eta^{\prime}}\frac{1}{\eta}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (53)
$=\frac{1}{\eta^{\prime}}\left(\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (54)
$=\frac{1}{\eta^{\prime}}\frac{\boldsymbol{p}_{j}}{\boldsymbol{q}_{j}}(\boldsymbol{n}_{j}^{\prime}\boldsymbol{v}_{j}^{\prime}-\boldsymbol{n}_{j}\boldsymbol{v}_{j}),$ (55)
due to all terms being equal except for $\boldsymbol{v}^{\prime}_{j}\neq\boldsymbol{v}_{j}$, $\boldsymbol{n}^{\prime}_{j}\neq\boldsymbol{n}_{j}$. ∎
State Expansion.
This case is the same as action backpropagation, where the updated index is a new summand at index $j=M+1$. The same equations apply with $n(s^{\prime,j})=0$.
Proof.
We append new terms to all vectors: $\boldsymbol{n}_{M+1}$, $\boldsymbol{p}_{M+1}$, $\boldsymbol{v}_{M+1}$, $\boldsymbol{q}_{M+1}$. Update to $\eta$:
$\eta^{\prime}-\eta$ (56)
$=\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}^{\prime}_{i}}{\boldsymbol{q}^{\prime}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}}{\boldsymbol{q}_{i}}$ (57)
$=\frac{\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}},$ (58)
due to all terms being equal except for index $M+1$. This corresponds to (51) with $\boldsymbol{n}_{M+1}=0$, after initializing $\boldsymbol{p}^{\prime}_{M+1}=\boldsymbol{p}_{M+1}$ and $\boldsymbol{q}^{\prime}_{M+1}=\boldsymbol{q}_{M+1}$ Similarly, for $\hat{V}_{f}(s,a)$:
$\hat{V}_{f}^{\prime}(s,a)-\frac{\eta}{\eta^{\prime}}\hat{V}_{f}(s,a)$ (59)
$=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\frac{\eta}{\eta^{\prime}}\frac{1}{\eta}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}$ (60)
$=\frac{1}{\eta^{\prime}}\left(\sum_{i=1}^{M+1}\frac{\boldsymbol{n}_{i}^{\prime}\boldsymbol{p}_{i}\boldsymbol{v}_{i}^{\prime}}{\boldsymbol{q}_{i}}-\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}\right)$ (61)
$=\frac{1}{\eta^{\prime}}\frac{\boldsymbol{p}^{\prime}_{M+1}}{\boldsymbol{q}^{\prime}_{M+1}}\boldsymbol{n}_{M+1}^{\prime}\boldsymbol{v}_{M+1}^{\prime},$ (62)
which again corresponds to (55) with $\boldsymbol{n}_{M+1}=0$. ∎
Action Update.
Let $a^{\prime}$ be the new action. Here, we have to calculate $\omega^{i}$ for all $s^{\prime,i}\in C(s,a)$ in $O(M)$ time. We update
$(\omega^{i})^{\prime}=p_{T}(s^{\prime,i}\mid s,a^{\prime})(p_{T}(s^{\prime,i}\mid s,a(s^{\prime,i})))^{-1},$ (63)
and recalculate (30) and (31) with the new ($\omega^{i})^{\prime}$, which are also $O(M)$ time operations. For the immediate reward, new rewards must be computed for the new triplets $(s,a^{\prime},s^{\prime,i})$, and the update is analogous.
Therefore, the log-form equations are:
$l(\eta^{\prime})=LSE(l(\boldsymbol{n})\oplus l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}))$ (64)
$\hat{V}_{f}^{\prime}(s,a)=\exp\left(LSE(l(\boldsymbol{p}^{\prime})\ominus l(\boldsymbol{q}),\boldsymbol{n}\odot\boldsymbol{v})-l(\eta^{\prime})\right)$ (65)
Proof.
This update corresponds to computing a new target likelihood vector $\boldsymbol{p}$. Therefore, the update equations follow from the definitions:
$\eta^{\prime}=\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}}{\boldsymbol{q}_{i}},$ (66)
$\hat{V}_{f}^{\prime}(s,a^{\prime})=\frac{1}{\eta^{\prime}}\sum_{i=1}^{M}\frac{\boldsymbol{n}_{i}\boldsymbol{p}_{i}^{\prime}\boldsymbol{v}_{i}}{\boldsymbol{q}_{i}}.$ (67)
Terminal State Backpropagation.
Let $s$ be a terminal node. Its value estimate is based only on rollouts. Hence, for a new rollout value $v^{\prime}$, we perform a running average update
$\hat{V}^{\prime}(s)=\hat{V}(s)+\frac{n^{\prime}(s)-n(s)}{n^{\prime}(s)_{+1}}(v^{\prime}-\hat{V}(s)),$ (68)
where $n^{\prime}(s)$ is $n(s)$ plus the number of rollouts (usually 1).
Proof.
This is a simple running average update for the mean estimator:
$\hat{V}(s)=\frac{1}{n(s)_{+1}}\sum_{i=1}^{n(s)_{+1}}v^{i},$ (69)
for all $v^{i}$ rollout values from the terminal node $s$. ∎
Non-Terminal State Backpropagation.
Let $(s,a^{j})$ be an updated node. Hence, we updated $n(s,a^{j})$ to $n^{\prime}(s,a^{j})$ and $\hat{Q}(s,a^{j})$ to $\hat{Q}^{\prime}(s,a^{j})$. We update $n^{\prime}(s)$ by (28) and perform
$\hat{V}^{\prime}(s)=(n^{\prime}(s))^{-1}\Big{(}n(s)\hat{V}(s)+n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})$
${}-n(s,a^{j})\hat{Q}(s,a^{j})\Big{)}.$ (71)
Proof.
Following the definition of the state value estimator (29), the updated estimator is:
$\hat{V}^{\prime}(s)=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i}).$ (72)
Therefore:
$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$ (73)
$=\sum_{C(s)}\frac{n^{\prime}(s,a^{i})}{n^{\prime}(s)}\hat{Q}^{\prime}(s,a^{i})-\frac{n(s)}{n^{\prime}(s)}\sum_{C(s)}\frac{n(s,a^{i})}{n(s)}\hat{Q}(s,a^{i})$ (74)
$=\frac{1}{n^{\prime}(s)}\sum_{C(s)}(n^{\prime}(s,a^{i})\hat{Q}^{\prime}(s,a^{i})-n(s,a^{i})\hat{Q}(s,a^{i})).$ (75)
All terms are equal except for the updated node $j$. Hence, we obtain
$\hat{V}^{\prime}(s)-\frac{n(s)}{n^{\prime}(s)}\hat{V}(s)$
$=\frac{1}{n^{\prime}(s)}\left(n^{\prime}(s,a^{j})\hat{Q}^{\prime}(s,a^{j})-n(s,a^{j})\hat{Q}(s,a^{j})\right).$ (76)
A.3 Theorem 1
We first highlight that the required ”well-behaved” conditions for the theorem are for changing the order of derivatives and integrals by the Leibniz integral rule. A sufficient condition is that the integrand and all its partial derivatives are continuous in all variables, which is satisfied when both $p_{T},r$ satisfy these conditions. Stronger versions of this theorem exist, such as only requiring that the partial derivatives exist almost everywhere. As this is not the focus of our discussion, we leave that for future work. Additionally, as stated in the main paper, it must be satisfied that $p_{T}(s^{\prime}\mid s,a)=0$ implies $p_{T}(s^{\prime}\mid s,a^{\prime})=0$.
Theorem 1.
For well-behaved $p_{T}$, $r$, the MDP action gradient of the action-value function is given by the expectation
$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})=\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}$
$(\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
$\cdot(r(s,a^{\prime},s^{\prime})+\gamma V_{t+1}^{\pi}(s^{\prime}))+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime}))\bigg{]}.$ (77)
Additionally, for well-behaved $p_{T}$, $p_{O}$, $r$, the POMDP action gradient of the action-value function can be computed by replacing posterior belief likelihood terms with propagated belief likelihood terms, i.e. $b_{t+1}^{-}$ instead of $b_{t+1}$ in all $p_{T}$ terms.
Proof.
We start with the MDP case.
$\nabla_{a^{\prime}}Q_{t}^{\pi}(s,a^{\prime})$ (78)
$=\nabla_{a^{\prime}}\mathbb{E}_{s{{}^{\prime}}\mid s,a}\bigg{[}\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))\bigg{]},$ (79)
Applying the Leibniz integral rule:
$=\mathbb{E}_{s{{}^{\prime}}\mid s,a}\bigg{[}\nabla_{a^{\prime}}\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}\cdot(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))$
$+\frac{p(s{{}^{\prime}}\mid s,a{{}^{\prime}})}{p(s{{}^{\prime}}\mid s,a)}\cdot\nabla_{a^{\prime}}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))\bigg{]}.$ (80)
For the first term, we use the log-gradient substitution:
$\nabla_{a^{\prime}}p(s{{}^{\prime}}\mid s,a{{}^{\prime}})=p(s{{}^{\prime}}\mid s,a{{}^{\prime}})\cdot\nabla_{a^{\prime}}\log p(s{{}^{\prime}}\mid s,a{{}^{\prime}}),$ (81)
and for the second term, since $V_{t+1}^{\pi}(s^{\prime})$ is not a function of $a^{\prime}$, we obtain:
$\nabla_{a^{\prime}}(r(s,a{{}^{\prime}},s{{}^{\prime}})+\gamma V_{t+1}^{\pi}(s{{}^{\prime}}))=\nabla_{a^{\prime}}r(s,a{{}^{\prime}},s{{}^{\prime}}).$ (82)
By gathering terms we obtain the result (77).
We now show the POMDP case. We assume that the same conditions are satisfied for $p_{T},p_{O},r$. We first show the propagated belief trick.
$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (83)
$=\int_{b^{\prime}}p(b^{\prime}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime},$ (84)
and we continue by marginalizing over the propagated belief $b^{-}$ and the observation $o$:
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{-},o,b^{\prime}\mid b,a)$
$(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}.$ (85)
By its definition, the propagated belief $b^{-}$ is the generated belief by conditioning on the history without the last observation. Therefore, using the chain rule and removing variables that are conditionally independent, we obtain:
[…]
$p(b^{-}\mid b,a)(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (86)
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(b^{\prime}\mid o,b^{-})p(o\mid b^{-})p(b^{-}\mid b,a)$
$(r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (87)
$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right].$ (88)
We introduce an importance ratio over the propagated belief to calculate the derivative for proposal action $a$ and target action $a^{\prime}$:
$Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))\bigg{]},$ (89)
where we denoted $p_{T}(b^{-}\mid b,a^{\prime})\triangleq p(b^{-}\mid b,a^{\prime})$ as the transition model for the belief MDP. Similarly to the derivation of the MDP case, we can apply the Leibniz integral rule to obtain the result:
$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a^{\prime})=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$(\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})$
$\cdot(r(b,a^{\prime},b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime}))+\nabla_{a^{\prime}}r(b,a^{\prime},b^{\prime}))\bigg{]}.$ (90)
A.4 Theorem 2
We highlight that as an extension of Theorem 1, it requires the same conditions. Additionally, for the simplification of the immediate reward expectation, we make use of Fubini’s theorem for changing integration orders. Sufficient conditions are that there exists some $R\in\mathbb{R}$ such that $\sup_{s,a,s^{\prime}}\lvert{r(s,a,s^{\prime})}\rvert<R$, and that $r,p_{T},p_{O}$ are all continuous.
Theorem 2.
For a state-reward POMDP, the action gradient of the action-value function is given by:
$\nabla_{a^{\prime}}Q_{t}^{\pi}(b,a)=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a^{\prime}}[\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})\cdot r(s,a^{\prime},s^{\prime})$
$+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})]+\gamma\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\bigg{]}.$ (91)
Proof.
The central part of this proof is the decomposition of the belief action-value function to an expectation over state trajectories of immediate rewards, and an expectation over belief trajectories of future values. Afterwards, we apply the Leibniz integral rule similarly to Theorem 1.
We start with the belief action-value function:
$Q_{t}^{\pi}(b,a)=\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})+\gamma V_{t+1}^{\pi}(b^{\prime})\right]$ (92)
$\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]+\gamma\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right].$ (93)
We expand the first expectation, using the chain rule and then Fubini’s theorem to cancel the expectation over the observation and posterior belief:
$\mathbb{E}_{b^{\prime}\mid b,a}\left[r(b,a,b^{\prime})\right]$ (94)
$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\left[\mathbb{E}_{s,s^{\prime}\mid b,a,b^{\prime}}[r(s,a,s^{\prime})]\right]$ (95)
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(s,s^{\prime},o,b^{-},b^{\prime}\mid b,a)$
$r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (96)
$=\int_{b^{-}}\int_{o}\int_{b^{\prime}}\int_{s}\int_{s^{\prime}}p(o,b^{\prime},b^{-}\mid s,s^{\prime},b,a)p(s,s^{\prime}\mid b,a)$
$r(s,a,s^{\prime})\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}$ (97)
$=\int_{s}\int_{s^{\prime}}p(s,s^{\prime}\mid b,a)r(s,a,s^{\prime})$
(∫$\cancel{\left(\int_{b^{-}}\int_{o}\int_{b^{\prime}}p(o,b^{\prime},b^{-}\mid s,s^{\prime},b,a)\mathop{}\!\mathrm{d}b^{\prime}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}b^{-}\right)}\mathop{}\!\mathrm{d}s^{\prime}\mathop{}\!\mathrm{d}s$ (98)
[…]
$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}[r(s,a,s^{\prime})].$ (100)
Now, we compute $\nabla_{a^{\prime}}$ for the terms obtained. The gradient of the immediate reward expectation, when the proposal action is $a$ and the target action is $a^{\prime}$, is given by (skipping some steps similar to Theorem 1):
$\nabla_{a^{\prime}}\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}r(s,a^{\prime},s^{\prime})\bigg{]}$ (101)
$=\mathbb{E}_{s\mid b}\mathbb{E}_{s^{\prime}\mid s,a}\bigg{[}\frac{p_{T}(s^{\prime}\mid s,a^{\prime})}{p_{T}(s^{\prime}\mid s,a)}\big{(}\nabla_{a^{\prime}}\log p_{T}(s^{\prime}\mid s,a^{\prime})$
$\cdot r(s,a^{\prime},s^{\prime})+\nabla_{a^{\prime}}r(s,a^{\prime},s^{\prime})\big{)}\bigg{]},$ (102)
and the importance ratio cancels out for $a=a^{\prime}$.
Next, the gradient of the future value expectation is simply:
$\nabla_{a^{\prime}}\mathbb{E}_{b^{\prime}\mid b,a}\left[V_{t+1}^{\pi}(b^{\prime})\right]$ (103)
$=\nabla_{a^{\prime}}\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}V_{t+1}^{\pi}(b^{\prime})\bigg{]}$ (104)
$=\mathbb{E}_{b^{-}\mid b,a}\mathbb{E}_{o\mid b^{-}}\mathbb{E}_{b^{\prime}\mid b^{-},o}\bigg{[}\frac{p_{T}(b^{-}\mid b,a^{\prime})}{p_{T}(b^{-}\mid b,a)}$
$\nabla_{a^{\prime}}\log p_{T}(b^{-}\mid b,a^{\prime})\cdot V_{t+1}^{\pi}(b^{\prime})\bigg{]}.$ (105)
Combining the two results, we obtain the final expression (91). ∎
Appendix B Experiments
B.1 AGMCTS Implementation Details
Optimizer
For the Adam optimizer, we used the hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$. The learning rate $\eta_{Adam}$ was determined by cross entropy optimization (CE), described later, in conjunction with the optimization over the other algorithm hyperparameters.
In our implementation we used an exponential decay on the step size after the Adam step size. The calculation for the update of the accumulated action is done as:
δAdam=
A
DAM
(a,gaq)
[…]
A
DAM
[…]
$\lambda=\max\{$0.999$^{T},$0.1$\}a^{acc}=a+\lambda\delta_{Adam},$ (107)
where $T$ is the number of gradient optimizations performed at this particular action node.
B.2 D-Continuous Light Dark
In this domain, $\mathcal{X}=\mathcal{Z}=\mathbb{R}^{D}$. The agent starts at a random position on the sphere of radius $r_{0}$ centered at the origin: $S^{D-1}_{r_{0}}$. The goal position is located at $x_{g}=(\boldsymbol{0}_{D-1},r_{g})$. A beacon is located to the side at $x_{b}=(r_{b},\boldsymbol{0}_{D-1})$. The constants we took are $r_{a}=1.5$, $r_{g}=r_{b}=2.5$, $r_{0}=0.5$. $T=0.2$ is the goal tolerance, and the POMDP terminates if $\lVert{x-x_{g}}\rVert<T$, or after $L=6$ time steps. We set $\gamma=0.99$.
The action space is $B_{r_{a}}(\boldsymbol{0}_{D})$. The transition model is a simple gaussian with added noise: $x^{\prime}\sim\mathcal{N}(x+a,\sigma_{T}^{2}\mathbf{I})$ with $\sigma_{T}=0.025$.
The observations are the agent’s relative position to the beacon, with a gaussian added noise that indreases with distance: $z\sim\mathcal{N}(x-x_{b},(\sigma_{Z}(\lVert{x-x_{b}}\rVert))^{2}\mathbf{I})$, with noise function $\sigma_{Z}(x)=\min\{\sigma_{Z}^{\max},k_{\sigma_{Z}}(x+x^{\alpha_{\sigma_{Z}}})\}$, for $\sigma_{Z}^{\max}=15$, $k_{\sigma_{Z}}=0.01$, $\alpha_{\sigma_{Z}}=8$.
The reward function is $r(x,a,x^{\prime})=R_{goal}\exp(-\frac{1}{2}{(\frac{d}{0.5T})}^{2})-R_{moat}\exp(-\frac{1}{2}{(\frac{d-5T}{T})}^{2})-R_{dist}d^{2}$, where $d\mathbin{=}\lVert{x-x_{g}}\rVert$. We took constants $R_{goal}\mathbin{=}10$, $R_{moat}\mathbin{=}2$, $R_{dist}\mathbin{=}0.02$. This reward function only depends on the posterior state. The agent needs to carefully target the goal as the reward decreases quickly from the center, and advancing at random towards the goal will incur a penalty due to the moat.
To simulate more difficult domains, we include a noisy rollout policy. The rollout heads directly to $x_{g}$, but the computed action has an added noise of $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$. In our experiments we used $\sigma_{r}=0.1$ for parameter tuning, and additionally evaluated for $\sigma_{r}=0.2,0.3$.
The implementation of D-Continuous Light Dark was written using POMDPs.jl [?]. The gradient function $\nabla_{a}r$ was equal 00 in our setting, and $\nabla_{a}\log p_{T}$ was calculated using Enzyme.jl [?]. Enzyme.jl provided extremely efficient gradients, achieving the execution times almost as close to the original functions. We did not care to optimize for making primal and dual calls, which might have allowed to save some computation time. However, it seems that it would have had a relatively minor effect as the heaviest computation was calculating the belief transition importance ratio, which due to it being a ratio of two products, cannot be MC approximated without introducing bias. Therefore, we iterated over all belief particles for its computation. Yet, in more complex problem where the dynamics are more complex, it might prove beneficial to compute gradients in the same pass as the transition likelihood wherever it can be done.
B.3 Evaluations
Parameter Notation Explanation
$c$ UCT exploration bonus
$k_{a}$ Action progressive widening (factor)
$\alpha_{a}$ Action progressive widening (power)
$k_{o}$ Observation progressive widening (factor)
$\alpha_{o}$ Observation progressive widening (power)
$\eta_{Adam}$ Adam step size (i.e. learning rate)
$T_{d_{a}}$ Action update distance threshold
Table 2: Hyperparameters optimized by CE in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions.
Parameter D POMCPOW PFT-DPW AGMCTS
$c$ 2 0.983 1.689 4.026
$k_{a}$ 2 0.350 7.332 8.346
$\alpha_{a}$ 2 0.834 0.473 0.515
$k_{o}$ 2 0.215 10.49 12.03
[…]
$c$ 3 1.024 2.429 5.212
$k_{a}$ 3 0.485 7.309 8.075
$\alpha_{a}$ 3 0.582 0.326 0.471
$k_{o}$ 3 0.744 11.27 15.20
$\alpha_{o}$ 3 0.226 0.195 0.317
[…]
$c$ 4 1.259 1.111 2.625
$k_{a}$ 4 0.360 9.309 8.043
$\alpha_{a}$ 4 0.559 0.343 0.495
$k_{o}$ 4 1.023 10.48 17.21
$\alpha_{o}$ 4 0.278 0.109 0.460
[…]
Table 3: Hyperparameters used in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions, rounded to $3$ significant digits after the decimal point.
We evaluated the algorithms on the D-Continuous Light Dark domain for $D=2,3,4$ dimensions. PFT-DPW and AGMCTS were given a budget of 500 simulations. The particle count during planning was set to be dependent on the dimension:
$J_{D=2}=256,\quad J_{D=3}=512,$
$J_{D=4}=1024.$ (108)
The particle filter used for inference between planning steps had a particle count $J^{\text{PF}}$ given by:
$J^{\text{PF}}_{D=2}=2048,\quad J^{\text{PF}}_{D=3}=4096,$
$J^{\text{PF}}_{D=4}=8192.$ (109)
POMCPOW was given a simulation budget according to the formula
$n_{D}^{sims,POMCPOW}=500\cdot$0.08$\cdot J_{D},$ (110)
which was empirically found to produce similar runtimes to PFT-DPW in the chosen domain.
In order to save computation time and not compute a partially observable rollout, the rollout computation for a new particle belief was done by drawing $s_{0}$ and 10 additional particles from the belief. The rollout actions were computed based on $s_{0}$, but the return was the mean of returns from the 10 particles. This mechanism allows to approximate a partially observable rollout without receiving any future observations. For POMCPOW, since every new node is expanded initially only with a single particle, the rollout was computed with that particle. This is a disadvantage for POMCPOW as it made the rollout returns less meaningful, however it seemd to have achieved better results than PFT-DPW despite that.
The algorithm hyperparameters were first determined by cross entropy optimization (CE) [?], largely following [?], using Gaussian distributions over the continuous DPW hyperparameters:
  1. 1.
    UCT exploration bonus $c$.
  2. 2.
    Action progressive widening parameters $k_{a}$ and $\alpha_{a}$.
  3. 3.
    Observation progressive widening parameters $k_{o}$ and $\alpha_{o}$.
For AGMCTS, we also optimized over the Adam learning rate $\eta_{Adam}$, and the action update distance threshold $T_{d_{a}}$. We searched for hyperparameters for each dimension while using rollout policy noise $\sigma_{r}=0.1$, and the results for $\sigma_{r}=0.2,0.3$ were obtained using the same hyperparameters of the same dimension. We used 150 parameter samples, 40 simulations for each parameter sample to determine its mean return, used 30 elite samples to fit the new distribution, with smoothing the new distribution’s parameters:
$\mu_{\text{new}}=\alpha_{\mu}\mu_{\text{new}}+(1-\alpha_{\mu})\mu_{\text{old}},$ (111)
$\Sigma_{\text{new}}=\alpha_{\Sigma}\Sigma_{\text{new}}+(1-\alpha_{\Sigma})\Sigma_{\text{old}},$ (112)
where we chose $\alpha_{\mu}=0.8,\alpha_{\Sigma}=0.5$.
We performed 50 CE iterations for each algorithm in each dimension. The starting mean for $D=2$ was the same for all algorithms, and for $D=3,4$ we used the mean of the best hyperparameters found for $D-1$. The hyperparameters chosen were those from the CE iteration achieving the maximal mean of mean returns over the 30 elite samples. The parameters used for each algorithm are summarized in Table 3. After choosing the parameters, we reported the mean return over 1000 scenarios, with the standard error of the mean. Table 4 summarizes the results with the individual mean runtimes for each algorithm in each scenario.
D $\sigma_{r}$ POMCPOW time (s) PFT-DPW time (s) AGMCTS time (s)
2 0.1 $\boldsymbol{6.62}\pm 0.07$ 0.114 $5.71\pm 0.08$ 0.126 $\boldsymbol{6.67}\pm 0.06$ 1.665
2 0.2 $\boldsymbol{6.23}\pm 0.07$ 0.122 $4.74,0.08$ 0.129 $5.96\pm 0.08$ 1.730
2 0.3 $\boldsymbol{5.61}\pm 0.08$ 0.129 $4.01\pm 0.09$ 0.129 $\boldsymbol{5.49}\pm 0.09$ 1.838
3 0.1 $3.42\pm 0.08$ 0.216 $3.63\pm 0.08$ 0.299 $\boldsymbol{5.47}\pm 0.07$ 2.442
3 0.2 $1.87\pm 0.08$ 0.240 $1.79\pm 0.08$ 0.296 $\boldsymbol{3.82}\pm 0.1$ 2.416
3 0.3 $1.06\pm 0.07$ 0.253 $0.29\pm 0.06$ 0.309 $\boldsymbol{1.58}\pm 0.1$ 2.389
4 0.1 $2.58\pm 0.07$ 0.548 $2.36\pm 0.07$ 0.683 $\boldsymbol{4.26},0.08$ 10.576
4 0.2 $0.53\pm 0.05$ 0.640 $0.2\pm 0.05$ 0.707 $\boldsymbol{1.53}\pm 0.09$ 8.303
4 0.3 $\boldsymbol{-0.36}\pm 0.03$ 0.660 $-0.86\pm 0.02$ 0.736 $-1.13\pm 0.04$ 5.778
Table 4: Performance results for the different algorithms at D-Continuous Light Dark for $D=2,3,4$ dimensions. AGMCTS and PFT-DPW were run for 500 simulations with particle counts given by (108), and POMCPOW for an approximately equivalent time to PFT-DPW according to (110). The rollout policy’s standard deviation was $\sigma_{r}=0.1$ during hyperparameter tuning, and then the algorithms were also tested for $\sigma_{r}=0.2,0.3$. The reported results are mean returns over 1000 scenarios $\pm$ standard error. AGMCTS achieved the best performance in most scenarios.